# Adaptation of Agentic AI

### v2

@@ 3 Overview of Adaptation Paradigms of Agentic AI
In this section, we provide an overview of the adaptation paradigms that form the analytical basis of this paper. Our objective is to establish a unified framework for categorizing existing studies on agentic AI systems according to what is adapted (the agent or the tool) and how the adaptation signal is obtained. We summarize these perspectives into four canonical paradigms, which together capture the major directions of adaptation explored in recent literature.
To facilitate a clear understanding of these paradigms, this section proceeds in three parts. We first introduce the mathematical notations that are used throughout this paper (§3.1). We then provide formal expressions for the four paradigms (§3.2). Finally, we present illustrative examples that help clarify how each paradigm operates and how they differ in adaptation mechanisms (§3.3).
@@ 3.1 Mathematical Notations
To ensure consistency in the subsequent formalization, here, we introduce the key mathematical notations used throughout this paper. We organize the notations into three conceptual categories that together define the adaptation process of agentic AI systems: the adaptation targets, the adaptation data sources, and the adaptation objectives.
@@
This category specifies the entities that undergo adaptation within an agentic AI system.
@@
  • Tool ($\mathcal{T}$): The set of external callable components that extend the agent’s capabilities beyond its internal parameters. Tools can include retrievers, planners, executors, simulators, or other computational modules. In this paper, we also categorize the memory module within $\mathcal{T}$, since memory can be viewed as a dynamic and updatable database that interacts with and learns from the agent’s outputs. Typically, the retrieval process for accessing stored information is performed through a dedicated retriever or search tool, which allows the agent to query and integrate relevant past knowledge into its reasoning process.
@@
This category describes the sources from which the adaptation signals are obtained.
@@
Having defined the adaptation targets and data sources, we next describe the objective that guides the adaptation process, which quantifies performance or alignment quality.
  • Objective Function $\mathcal{O}(\cdot)$: The objective function optimized during adaptation, which evaluates how effectively the agent–tool system performs according to the designated evaluation protocol. For example, the objective for offline data $\mathcal{D}$ may correspond to supervised or imitation learning losses such as supervised fine-tuning (SFT) or behavior cloning. When adaptation relies on interactions with the environment $\mathcal{E}$, the objective is typically defined by outcome-based metrics such as task success rate.
These notations provide a unified foundation for expressing how adaptation operates at both the agent and tool levels, which we formalize in the subsequent subsection.
@@ 3.2 Four Adaptation Paradigms of Agentic AI
Building upon the mathematical notations introduced earlier, we now present the four adaptation paradigms proposed in this paper, which together form a unified framework for classifying existing approaches to agentic AI adaptation. In this framework, adaptation is first categorized by the optimization target, namely the agent or the tool. For agent adaptation, we further differentiate paradigms based on the type of optimization signal used, which may originate from tool-execution feedback (A1) or from evaluations of the agent’s own final output (A2). For tool adaptation, the distinction instead concerns whether the adaptation process involves the agent, where tools may be optimized independently of any agent (T1) or adapted under the supervision of a fixed agent (T2). Taken together, these considerations give rise to four paradigms, A1, A2, T1, and T2, which collectively characterize the principal modes of adaptation explored in agentic AI research.
@@ 3.2.1 A1: Tool Execution Signaled Agent Adaptation
This paradigm focuses on improving the agent $\mathcal{A}$ through feedback signals derived from the execution results of external tools $\mathcal{T}$. It captures scenarios where the agent interacts with tools in a verifiable manner, allowing the tool outcomes to serve as a measurable basis for optimization.
Agent–Tool Interaction Process.
The agent receives an input $x$ (e.g., a user query or task description) and generates a structured tool call or action $a=\mathcal{A}(x)$, which may include the tool name, arguments, and calling context. The tool set $\mathcal{T}$ then executes this call to produce a result $y=\mathcal{T}(a)$. The pair $(a,y)$ represents a single agent–tool interaction, and the overall process can be summarized as
@@
This pipeline captures how the agent leverages tools to complete tasks. For simplicity and without loss of generality, we describe the interaction using a single tool invocation; multi-turn tool use follows as a direct extension of the formulation.
@@ 3.2.2 A2: Agent Output Signaled Agent Adaptation
Unlike the A1 paradigm, where the adaptation signal is derived from tool-execution outcomes, the A2 paradigm obtains its optimization signal from the agent’s own final output. For simplicity and without loss of generality, the following description focuses on a single-turn interaction; multi-turn tool use, where the agent invokes multiple tool calls and integrates multiple intermediate results, extends naturally from the formulation.
Agent–Tool Interaction Process.
@@
  • Supervised Fine-Tuning (SFT). Let $\mathcal{D}_{\text{ans}}=\{(x,y,a^{*},o^{*})\}$ denote a dataset consisting of the input $x$, optional intermediate tool outputs $y$, the reference tool call $a^{*}$, and the corresponding target final output $o^{*}$. A key characteristic of the A2 paradigm is that its adaptation signal comes solely from the final agent output. However, supervising only the final output $o^{*}$, i.e., optimizing
    $\mathcal{A}^{*}=\arg\min_{\mathcal{A}}\,\mathbb{E}_{(x,y,a^{*},o^{*})\sim\mathcal{D}_{\text{ans}}}\Big[\ell\big(\mathcal{A}(x,a^{*},y),o^{*}\big)\Big]\;\equiv\;\arg\max_{\mathcal{A}}\,\mathbb{E}_{(x,y,a^{*},o^{*})}\Big[\log p_{\mathcal{A}}(o^{*}|x,a^{*},y)\Big],$ (5)
    is insufficient for learning tool-use behavior: the agent could improve its final-answer likelihood without ever invoking tools, since the supervision provides no incentive to produce the correct tool call. Therefore, for A2-style SFT to effectively support tool-using agents, it must combine final-output supervision with tool-call supervision, effectively integrating A2-style supervision with the A1-style imitation of tool-use trajectories. The supervised objective then becomes:
    $\mathcal{A}^{*}=\arg\min_{\mathcal{A}}\,\mathbb{E}_{(x,y,a^{*},o^{*})\sim\mathcal{D}_{\text{ans}}}\Big[\ell\big(\mathcal{A}(x),a^{*}\big)+\ell\big(\mathcal{A}(x,a^{*},y),o^{*}\big)\Big],$
    which is equivalently written as:
    $\arg\max_{\mathcal{A}}\,\mathbb{E}_{(x,y,a^{*},o^{*})}\Big[\log p_{\mathcal{A}}(a^{*}|x)+\log p_{\mathcal{A}}(o^{*}|x,a^{*},y)\Big].$ (6)
    Here, $\ell$ denotes the cross-entropy loss used in next-token prediction for LLMs. The first term teaches the agent to make correct tool calls (A1-style imitation), while the second term supervises the agent’s final answer generation (A2-style imitation). This formulation naturally includes the special case where no tools are invoked, in which case $a^{*}$ and $y$ are empty.
@@ 3.2.3 T1: Agent-Agnostic Tool Adaptation
In the T1 paradigm, the agent $\mathcal{A}$ is kept fixed, and adaptation is applied to only the external tool set $\mathcal{T}$. This setting arises naturally when the agent is a powerful and robust closed-source API (such as GPT, Claude, or Gemini) that usually cannot be fine-tuned, or when the goal is to enhance the fixed agent by training specialized tools to complement the frozen agent, such as retrievers, rerankers, planners, simulators, or additional foundation models. In this sense, a “tool” in T1 primarily refers to a trainable model, regardless of whether it is a traditional machine-learning model or a large-scale foundation model.
@@ 3.2.4 T2: Agent-Supervised Tool Adaptation
In the T2 paradigm, tool adaptation is guided by the frozen agent $\mathcal{A}$. Unlike T1, where tools are trained independently of any agent, T2 explicitly aims to adapt or construct a tool that complements the fixed agent and enhances its overall capability. This setting reflects a practical motivation: when the main agent is a powerful closed-source foundation model, it is often preferable to train auxiliary tools around it rather than modifying the agent itself.
Agent–Tool Interaction Process.
Without loss of generality, we describe the interaction using a single-turn example, noting that multi-turn processes can be extended naturally from this. The agent receives input $x$ and produces a tool call $a=\mathcal{A}(x)$. The tool $\mathcal{T}$ executes this call to return a result $y=\mathcal{T}(a)$, and the agent integrates $(x,a,y)$ or $(x,y)$ to produce the final output $o$:
@@
The tool is optimized to improve the performance of the fixed agent–tool system:
@@
where $\mathcal{O}_{\text{agent}}$ evaluates how effectively the agent performs when equipped with tool $\mathcal{T}$. This objective emphasizes that T2 adapts the tool specifically to the needs of the given agent. Tool adaptation in T2 generally takes two forms:
@@
Memory as a Special Case of T2.
In this paper, the memory module storing long-term memory is treated as a tool within the T2 paradigm. Although memory serves a conceptual role distinct from conventional executable tools, its update mechanism aligns precisely with the agent-driven tool adaptation view. During interaction, the frozen agent produces a final output $o$ that reflects its reasoning over both the input and retrieved memory contents. This final output $o$ is then used to update the memory module through a fixed or learnable write function: $\mathcal{M}\leftarrow\text{Update}(\mathcal{M},\,o)$, where $\mathcal{M}$ denotes the memory store. This process corresponds directly to T2: the agent remains fixed, the adaptation signal originates from the agent’s own output, and the memory module is optimized to better support future agent reasoning.
Thus, adaptive memory systems naturally fall under the T2 paradigm: the tool $\mathcal{T}$ being optimized is the memory module, and the supervision signal arises entirely from the behavior of a fixed agent $\mathcal{A}$ interacting with and benefiting from that memory.
@@ 3.3 Illustrative Examples
To make the above adaptation paradigms more concrete, this section provides illustrative examples drawn from two representative application settings: retrieval-augmented generation (RAG) and code-execution-based tasks. These two settings are chosen because they highlight the central role of tool use in agentic AI systems, while exhibiting distinct tool–agent interaction patterns and evaluation protocols.
For each application, we present a pair of examples that correspond to the A1 and A2 paradigms. The paired examples share the same form of tool-call action, i.e., document retrieval or executing code, respectively. This allows us to clearly contrast tool-feedback–based agent adaptation (A1) with agent-output–based agent adaptation (A2). Finally, we provide a T2 example in the RAG setting. By examining these examples side by side, readers can develop an intuitive understanding of how these paradigms differ in objectives, update signals, and learning dynamics.
@@ 3.3.1 Agent Adaptation Examples Across Two Applications
We now illustrate the A1 and A2 paradigms through two representative tool-use settings: retrieval-augmented generation (RAG) and code-execution-based question answering. For each application, we begin by describing the underlying problem setup, followed by a pair of examples that instantiate A1 and A2 under the same form of tool-call action. This allows a clean comparison between adaptation driven by tool-execution feedback (A1) and adaptation driven by agent final outputs (A2).
@@
  • A1 example: DeepSeek-R1 (code) [24]. During reinforcement learning, DeepSeek-R1 generates code that is executed inside a sandbox. The execution output, such as test-case pass rate or numerical correctness, is used directly as the reward for policy optimization. Since adaptation is based entirely on the tool’s execution result, this example fits the A1 paradigm.
@@ 3.3.2 Tool Adaptation Examples in the RAG Setting
In many practical systems, the central agent is instantiated as a powerful closed-source API model (such as GPT-, Claude-, or Gemini-style models) that already exhibits strong performance and robustness across a wide range of tasks. Training from an open-source model to match this level of robustness is extremely challenging: the data quality must be carefully curated, scaling laws suggest that much larger models and much more data are required for competitive performance, and such models demand substantial training infrastructure. As a result, a convenient and often more feasible strategy is to treat the closed-source API model as a fixed agent and instead adapt auxiliary tools around it. In the RAG setting, this motivates tool adaptation for components such as retrievers, which can be optimized to complement the fixed agent.
@@
(1) Classic Dense Retrievers. Under the T1 paradigm, tools are trained independently of any specific agent and can be plugged into frozen LLM agents without further co-adaptation. A canonical example is a standard dense retriever, such as a bi-encoder trained with contrastive learning to map queries and documents into a shared embedding space for vector similarity search. Once trained, such a retriever can be used as a standalone retrieval tool: given a tool call action $a$ (a retrieval query), the dense retriever returns a ranked document set $y=\mathcal{T}(a)$ optimized for recall or semantic relevance. A closed-source agent (e.g., GPT-, Claude-, or Gemini-style) can then consume $y$ to perform downstream reasoning and answer synthesis, despite the agent itself never participating in the retriever’s training. (2) Learned Subagents as Agent-Agnostic Tools. Beyond classic dense retrievers, once agent adaptation has produced strong retrieval-oriented models, these learned models can themselves be reused as tools under the T1 paradigm. For example, a model trained in the DeepRetrieval [21] style can be deployed purely as a high-quality subagent that rewrites queries for improved retrieval over specific document databases. Given a tool call action $a$ (a retrieval query), the subagent returns a reformulated query or a curated document set $y=\mathcal{T}(a)$ with enhanced recall or relevance, which is then consumed by a fixed closed-source agent that performs the final reasoning and answer synthesis.
T2 examples:
@@
With these illustrative examples in place, we next move on to systematically review the literature associated with each of the four adaptation paradigms.
@@ 4 Agent Adaptation
Agent adaptation refers to the mechanisms through which agents refine their behavior and decision-making capabilities based on feedback from their interactions with tools, environments, or their own outputs. This process is pivotal for enhancing the autonomy, reasoning, and generalization abilities of agents across diverse tasks. Broadly, agent adaptation can be categorized into two paradigms: A1, which leverages tool execution results as feedback signals, and A2, which focuses on evaluating the agent’s own outputs.
Formally, let $\mathcal{A}$ denote an agent, parameterized by its internal configuration or policy (which includes prompt templates or model weights), and let $\mathcal{T}$ represent the set of tools accessible to the agent. The agent’s performance under a given configuration is evaluated by an objective function $\mathcal{O}(\cdot)$, which provides feedback based on either tool performance or agent output quality. Accordingly, the two adaptation paradigms can be formalized as optimization objectives:
@@
where $\mathcal{O}_{\text{tool}}$ quantifies the correctness or utility of outcomes derived from tool execution, such as successful code compilation or retrieval precision, and $\mathcal{O}_{\text{agent}}$ measures the quality of the agent’s generated outputs, including reasoning validity, factual accuracy, or alignment with human preferences. Here, $\mathcal{A}^{*}$ denotes the optimized agent configuration that maximizes the corresponding feedback objective.

### v3

@@ 3 Overview of Adaptation Paradigms of Agentic AI
We present the adaptation paradigms that form the analytical basis of this paper. We categorize existing studies on agentic AI systems according to what is adapted (the agent or the tool) and how the adaptation signal is obtained, yielding four canonical paradigms that capture the major directions of adaptation in recent literature.
@@ 3.1 Mathematical Notations
We introduce the key mathematical notations used throughout this paper, organized into three categories: adaptation targets, adaptation data sources, and adaptation objectives.
@@
The entities that undergo adaptation within an agentic AI system are as follows.
@@
  • Tool ($\mathcal{T}$): The set of external callable components that extend the agent’s capabilities beyond its internal parameters. Tools can include retrievers, planners, executors, simulators, or other computational modules. In this paper, we treat external memory modules as components of $\mathcal{T}$ when they function as dynamic, updatable stores that interact with and learn from the agent’s outputs. The retrieval process for accessing stored information is typically performed through a dedicated retriever or search tool, which allows the agent to query and integrate relevant past knowledge into its reasoning process. We note that parametric and hybrid memory forms may blur the boundary between $\mathcal{T}$ and $\mathcal{A}$; see the discussion below (§3.2).
@@
The sources from which adaptation signals are obtained include the following.
@@
The objective that guides the adaptation process quantifies performance or alignment quality.
  • Objective Function $\mathcal{O}(\cdot)$: The objective function optimized during adaptation, which evaluates how effectively the agent-tool system performs according to the designated evaluation protocol. For example, the objective for offline data $\mathcal{D}$ may correspond to supervised or imitation learning losses such as supervised fine-tuning (SFT) or behavior cloning. When adaptation relies on interactions with the environment $\mathcal{E}$, the objective is typically defined by outcome-based metrics such as task success rate.
@@ 3.2 Four Adaptation Paradigms of Agentic AI
Building on these notations, we present the four adaptation paradigms proposed in this paper. Adaptation is first categorized by the optimization target: the agent or the tool. For agent adaptation, paradigms are further differentiated by the optimization signal: tool-execution feedback (A1) or evaluations of the agent’s final output (A2). For tool adaptation, the distinction concerns whether tools are optimized independently of any agent (T1) or adapted under the supervision of a fixed agent (T2).
@@ 3.2.1 A1: Tool Execution Signaled Agent Adaptation
In the A1 paradigm, the agent $\mathcal{A}$ is improved through feedback signals derived from the execution results of external tools $\mathcal{T}$, capturing scenarios where tool outcomes serve as a measurable basis for optimization.
Agent-Tool Interaction Process.
The agent receives an input $x$ (e.g., a user query or task description) and generates a structured tool call or action $a=\mathcal{A}(x)$, which may include the tool name, arguments, and calling context. The tool set $\mathcal{T}$ then executes this call to produce a result $y=\mathcal{T}(a)$. The pair $(a,y)$ represents a single agent-tool interaction, and the overall process can be summarized as
@@
The pipeline captures how the agent uses tools to complete tasks. For simplicity and without loss of generality, we describe the interaction using a single tool invocation; multi-turn tool use follows as a direct extension of the formulation.
@@ 3.2.2 A2: Agent Output Signaled Agent Adaptation
Unlike the A1 paradigm, where the adaptation signal is derived from tool-execution outcomes, the A2 paradigm obtains its optimization signal from the agent’s own final output. For simplicity, the following description focuses on a single-turn interaction; the multi-turn case extends naturally.
Agent-Tool Interaction Process.
@@
  • Supervised Fine-Tuning (SFT). Let $\mathcal{D}_{\text{ans}}=\{(x,y,a^{*},o^{*})\}$ denote a dataset of inputs, optional tool outputs, reference tool calls, and target final outputs. Supervising only the final output $o^{*}$ is insufficient for learning tool-use behavior, because the agent could improve final-answer likelihood without ever invoking tools. Effective A2-style SFT therefore combines final-output supervision with A1-style tool-call imitation:
    $\mathcal{A}^{*}=\arg\max_{\mathcal{A}}\,\mathbb{E}_{(x,y,a^{*},o^{*})}\Big[\underbrace{\log p_{\mathcal{A}}(a^{*}|x)}_{\text{tool-call (A1-style)}}+\underbrace{\log p_{\mathcal{A}}(o^{*}|x,a^{*},y)}_{\text{final answer (A2-style)}}\Big].$ (5)
    When no tools are invoked, $a^{*}$ and $y$ are empty and the objective reduces to standard answer-level SFT.
@@ 3.2.3 T1: Agent-Agnostic Tool Adaptation
In the T1 paradigm, the agent $\mathcal{A}$ is kept fixed, and adaptation is applied to only the external tool set $\mathcal{T}$. The setting arises when the agent is a closed-source API (such as GPT, Claude, or Gemini) that cannot be fine-tuned, or when the goal is to augment a fixed agent by training specialized tools such as retrievers, rerankers, planners, simulators, or additional foundation models. In this sense, a “tool” in T1 primarily refers to a trainable model, regardless of whether it is a traditional machine-learning model or a large-scale foundation model.
@@ 3.2.4 T2: Agent-Supervised Tool Adaptation
In the T2 paradigm, tool adaptation is guided by the frozen agent $\mathcal{A}$. Unlike T1, where tools are trained independently, T2 adapts or constructs tools that complement the fixed agent and improve its overall capability. When the main agent is a closed-source foundation model, training auxiliary tools around it is often preferable to modifying the agent itself.
Agent-Tool Interaction Process.
As before, we describe a single-turn interaction; the multi-turn case extends naturally. The agent receives input $x$ and produces a tool call $a=\mathcal{A}(x)$. The tool $\mathcal{T}$ executes this call to return a result $y=\mathcal{T}(a)$, and the agent integrates $(x,a,y)$ or $(x,y)$ to produce the final output $o$:
@@
The tool is optimized to improve the performance of the fixed agent-tool system:
@@
where $\mathcal{O}_{\text{agent}}$ evaluates how effectively the agent performs when equipped with tool $\mathcal{T}$. The objective emphasizes that T2 adapts the tool specifically to the needs of the given agent. Tool adaptation in T2 generally takes two forms:
@@
Memory and the Adaptation Paradigms.
In this paper, the memory module is treated as a tool within $\mathcal{T}$. However, not all memory systems map cleanly onto a single paradigm; the appropriate classification depends on the memory’s form and update mechanism:
  • External adaptive memory (predominantly T2). When a frozen agent produces outputs that are used to update an external memory store (e.g., episodic buffers, reflective databases, skill libraries) through a fixed or learnable write function $\mathcal{M}\leftarrow\text{Update}(\mathcal{M},\,o)$, the process aligns with T2: the agent remains fixed, the adaptation signal originates from the agent’s own output, and the memory module evolves to better support future reasoning.
  • Pre-trained or plug-in memory modules (predominantly T1). Memory components that are trained independently of any specific agent, such as pre-trained dense retrievers, static knowledge bases, or off-the-shelf embedding indices, function as agent-agnostic tools under T1.
  • Parametric and hybrid memory (boundary cases). Memory encoded within model parameters (e.g., LoRA adapters for knowledge injection, differentiable memory modules like Titans) or hybrid architectures that combine parametric and external storage (e.g., Memory3) blur the boundary between tool adaptation and agent adaptation. When the memory update requires gradient-based changes to the agent’s own parameters, the method is better classified under A1 or A2; when only an auxiliary parameter set is updated while the core agent remains frozen, the method occupies the T1/T2 boundary.
We adopt the convention that the dominant form of memory adaptation determines the paradigm label. For most external, non-parametric memory systems discussed in this survey, the T2 label applies: the frozen agent’s outputs supervise the evolution of the memory module. We flag boundary cases explicitly in §5 to avoid flattening the important architectural differences among memory forms.
@@ 3.3 Illustrative Examples
We illustrate the adaptation paradigms through two representative settings: retrieval-augmented generation (RAG) and code-execution-based tasks. These settings highlight the central role of tool use in agentic AI while exhibiting distinct interaction patterns and evaluation protocols.
For each application, we present paired A1 and A2 examples sharing the same tool-call action (document retrieval or code execution), allowing a direct contrast between tool-feedback-based adaptation (A1) and agent-output-based adaptation (A2). We also provide a T2 example in the RAG setting.
@@ 3.3.1 Agent Adaptation Examples Across Two Applications
We illustrate the A1 and A2 paradigms through two tool-use settings: retrieval-augmented generation (RAG) and code-execution-based question answering. For each, we describe the problem setup and then present paired examples that instantiate A1 and A2 under the same tool-call action.
@@
Contrast. DeepRetrieval (A1) directly optimizes retrieval quality but provides no gradient signal for answer synthesis; Search-R1 (A2) optimizes end-to-end answer correctness but must discover good retrieval strategies as a side effect, making credit assignment harder.
@@
  • A1 example: DeepSeek-R1 (code) [25]. During reinforcement learning, DeepSeek-R1 generates code that is executed inside a sandbox (when the reward is derived solely from test-case pass rates, this is A1). The execution output, such as test-case pass rate or numerical correctness, is used directly as the reward for policy optimization. Since adaptation is based entirely on the tool’s execution result, this example fits the A1 paradigm.
@@
Contrast. DeepSeek-R1’s A1 formulation ties the reward to test-case pass rates, giving the agent a precise signal per code invocation. ReTool’s A2 formulation, by contrast, rewards only the final answer, so the model must learn when to invoke the code sandbox—a richer but noisier learning problem.
@@ 3.3.2 Tool Adaptation Examples in the RAG Setting
In many practical systems, the central agent is a closed-source API model (e.g., GPT, Claude, or Gemini) that cannot be fine-tuned. Training an open-source model to match such performance is challenging due to data curation requirements, scaling laws, and infrastructure demands. A more feasible strategy is to treat the closed-source model as a fixed agent and adapt auxiliary tools around it. In the RAG setting, this motivates tool adaptation for components such as retrievers.
@@
(1) Classic Dense Retrievers. A standard dense retriever (bi-encoder trained with contrastive learning) is the canonical T1 tool: given a query $a$, it returns documents $y=\mathcal{T}(a)$ optimized for recall, and any frozen agent can consume $y$ for downstream reasoning without having participated in the retriever’s training. (2) Learned Subagents as Agent-Agnostic Tools. Beyond classic dense retrievers, once agent adaptation has produced strong retrieval-oriented models, these learned models can themselves be reused as tools under the T1 paradigm. For example, a model trained in the DeepRetrieval [22] style can be deployed purely as a high-quality subagent that rewrites queries for improved retrieval over specific document databases. Given a tool call action $a$ (a retrieval query), the subagent returns a reformulated query or a curated document set $y=\mathcal{T}(a)$ with higher recall or relevance, which is then consumed by a fixed closed-source agent that performs the final reasoning and answer synthesis.
T2 examples.
@@ 4 Agent Adaptation
Agent adaptation refers to the mechanisms through which agents refine their behavior based on feedback from tools, environments, or their own outputs. A central design choice is where the feedback signal originates, which determines both what the agent can learn and what failure modes remain invisible. When the signal comes from tool execution (A1), the agent receives dense, causally grounded feedback tied to specific actions—but this signal is blind to whether the agent’s overall reasoning strategy is sound. When the signal comes from evaluating the agent’s own outputs (A2), the agent can optimize holistic task success—but the sparser, episode-level reward makes credit assignment harder and training less data-efficient. This A1/A2 tension between mechanistic precision and strategic flexibility structures the design space reviewed in this section.
Formally, let $\mathcal{A}$ denote an agent parameterized by its internal configuration or policy (prompt templates or model weights), and let $\mathcal{T}$ represent the set of accessible tools. The two paradigms correspond to distinct optimization objectives:
@@
where $\mathcal{O}_{\text{tool}}$ scores the correctness or utility of tool execution outcomes (e.g., code compilation success, retrieval precision), and $\mathcal{O}_{\text{agent}}$ scores the quality of the agent’s generated outputs (reasoning validity, factual accuracy, or preference alignment). The distinction has practical consequences: $\mathcal{O}_{\text{tool}}$ yields per-action credit but is silent about reasoning quality, whereas $\mathcal{O}_{\text{agent}}$ captures end-to-end performance but conflates tool-use skill with reasoning skill, complicating diagnosis when either degrades.